The length of a sonometer is 110 cm, then where should be 2 bridges be placed from A to divide the wire in 3 segments whose fundamental frequincies are in 1:2:3 ratio ?

The length of  a sonometer is 110 cm, then where should be 2 bridges be placed from A to divide the wire in 3 segments whose fundamental frequincies are in 1:2:3 ratio ? 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
as
n1 :n2 : n3 = 1: 2: 3
hence
l1 : l2 : l3 = 1 : ½ : 1/3
sum of the ratio = 6+ 3 + 2 = 11
hence l1 = 110 * 6 /11 = 60 cm
l2 = 110 * 3 / 11 = 30 cm
thus the wedges should be placed from A at 60cm and 90 cm.